OpenStudy (anonymous):

Khang and Hector live 32 miles apart in southeastern Missouri. They decide to bicycle towards each other and meet somewhere in between. Hector's rate of speed is 60% of Khang's. They start out at the same time and meet 2 hours later. Find Hector's rate of speed.

OpenStudy (anonymous):



OpenStudy (anonymous):

Khang will travel dk=V*t=v*2 miles in two hours. Hector will travel dh=V*t=0.6v*2 miles in two hours. Finally, dk+dh must equal 32 miles. You can find v then :)